# Diversity as a Reward: Fine-Tuning LLMs on a Mixture of Domain-Undetermined Data

### v1

@@ Appendix D More Experimental Results
Figure 9: The t-SNE visualization of embeddings for data samples with different distributions on Qwen2.5-7B. (a) The data pool of all 40,000 samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids based on Inter-Diversity, (d) Distribution of data closest to other domain centroids based on Inter-Diversity, (e) Distribution of data closest to its own domain centroid based on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid based on Inter-Diversity.
D.2 The Visualization of Inter- & Intra-Diversity Distribution Data
@@
D.3 Similarity in Generated Centroid Quantities
To assess the impact of the amount of generated data on the accuracy of Domain Semantic Centroids, we select quantities of 10 and 30 samples for generation. Subsequently, we calculate the semantic cosine similarity between these two sets, with results presented in Table 5 (using Qwen2-7B and Llama3.1-8B as examples). The observed differences between varying data quantities were not significant, indicating that the generalization of the generated data is sufficient and further increasing the data quantity does not significantly enhance the accuracy of the centroids.
Table 5: Semantic cosine similarity between generated samples of size 10 and 30 across different domains.
Similarity of 10 & 30 Common Sense Reasoning Mathematics Coding
Qwen2-7B 0.9686 0.9866 0.9919 0.9881
Llama3.1-8B 0.9851 0.9795 0.9685 0.9895
D.4 Similarity in Generated Centroid Domains
D.5 DaaR Layer Selection Protocol
To determine an appropriate layer for embedding DaaR into a large language model (LLM) that balances performance and computational cost, we first evaluate the capacity of different layers for domain awareness. Using Qwen2-7B as an example, we conduct t-SNE visualizations of Layer-5, Layer-10, Layer-15, Layer-20, Layer-25, and the last hidden layer, based on the average token vectors, as shown in Fig. 13.
Our observations indicate that even at the early Layer-5, the model begins to exhibit classification capabilities across different datasets. This prompted us to focus on the initial five layers to enhance learning outcomes. Considering that data clustering is inherently influenced by the embedding layer, a natural approach was to directly connect to the embedding layer for training. However, the training results showed persistently high training loss and a validation accuracy of only about 0.6, indicating suboptimal learning performance. Furthermore, related studies indicate that LLMs can learn semantic understanding and perception in the early layers (Wei et al., 2022), consequently, we select Layer-3 for DaaR.
Figure 14: Training loss and validation process of Embedding Layer on Qwen2-7B.
D.6 Comprehensive Training Dynamics
@@
D.7 Complete Validation Results of Inter-Diversity and Intra-Diversity
@@
Table 6: Validation Results of Inter-Diversity and Intra-Diversity on Llama3.1-8B across various benchmarks.
@@
Table 7: Validation Results of Inter-Diversity and Intra-Diversity on Qwen2-7B across various benchmarks.
@@
Table 8: Validation Results of Inter-Diversity and Intra-Diversity on Qwen2.5-7B across various benchmarks.
@@
D.8 Complete Results of DaaR with Baselines
@@
Alpagasus (Chen et al., 2024d) 21.57 64.37 74.87 55.20 17.65 4.60 16.16 36.34
Instag-C (Lu et al., 2024) 18.12 64.96 74.01 55.70 15.50 4.80 37.81 38.70
Instag-D (Lu et al., 2024) 21.94 64.69 74.87 54.80 12.80 4.10 9.76 34.71
SuperFilter (Li et al., 2024a) 22.95 64.99 76.39 57.60 6.05 2.60 40.55 38.73
Deita-C (Liu et al., 2024b) 15.58 64.97 74.21 55.00 13.05 4.60 34.46 37.41
Deita-Q (Liu et al., 2024b) 19.57 64.22 75.15 54.00 7.20 4.20 28.35 36.10
Deita-D (Liu et al., 2024b) 20.97 63.32 75.10 54.90 7.00 4.00 31.71 36.71
@@
Alpagasus (Chen et al., 2024d) 13.67 57.94 73.04 73.90 32.30 51.40 63.41 52.24
Instag-C (Lu et al., 2024) 9.51 58.50 73.06 74.70 35.35 51.90 64.70 52.53
Instag-D (Lu et al., 2024) 12.87 57.48 72.80 74.40 33.75 51.80 64.02 52.45
SuperFilter (Li et al., 2024a) 19.16 58.98 72.99 73.70 30.10 52.40 58.85 52.31
Deita-C (Liu et al., 2024b) 8.94 58.07 73.06 73.90 35.55 52.90 62.20 52.09
Deita-Q (Liu et al., 2024b) 14.06 59.07 73.16 75.80 35.50 23.00 58.24 48.40
Deita-D (Liu et al., 2024b) 16.41 57.80 72.70 76.10 29.05 52.40 64.63 52.73
@@
Alpagasus (Chen et al., 2024d) 10.40 57.87 72.92 77.20 18.75 61.80 65.55 52.07
Instag-C (Lu et al., 2024) 10.81 58.45 73.27 76.00 13.30 61.80 68.29 51.70
Instag-D (Lu et al., 2024) 11.08 58.40 72.79 76.40 16.40 62.90 70.43 52.63
SuperFilter (Li et al., 2024a) 13.54 58.51 72.89 79.30 11.35 39.50 65.25 48.62
Deita-C (Liu et al., 2024b) 10.50 58.17 73.14 74.60 16.60 62.00 72.26 52.47
Deita-Q (Liu et al., 2024b) 11.24 57.83 72.97 78.50 12.95 38.10 67.68 48.47
Deita-D (Liu et al., 2024b) 10.48 57.81 73.05 77.20 15.25 52.90 69.21 50.84

### v2

@@ F.2 Validation on Qwen3-8B
Figure 8: The t-SNE visualization of embeddings for data samples with different distributions on Qwen3-8B. (a) The data pool of all 40K samples, (b) Distribution of data farthest from other domain centroids on Inter-Diversity, (c) Distribution of data closest to other domain centroids on Inter-Diversity, (d) Distribution of data closest to its own domain centroid on Inter-Diversity, (e) Distribution of data farthest from its own domain centroid on Inter-Diversity.
Subsequently, we processed the curated dataset through inter-diversity and intra-diversity filters, followed by training Qwen3 on artificially constructed data distributions. As evidenced in Table 8, the experimental outcomes reveal that while the performance patterns diverge from those observed in Llama 3.1, Qwen2, and Qwen2.5, varying diversity levels demonstrably influence comprehensive model performance. Notably, the absolute value of the comprehensive performance of Qwen3 is even inferior to that of Qwen2.5 and Qwen2. We posit that this phenomenon is primarily due to the incomplete support of the evaluation platform OpenCompass (as of the submission date), where the distinctive thinking and reasoning capabilities of Qwen3 have not been fully leveraged. However, the relative comparison of different processes within the same data pool remains highly informative.
Table 8: Validation results of Inter-Diversity and Intra-Diversity on Qwen3-8B across benchmarks.
Qwen3-8B Common Sense Reasoning Mathematics Coding Avg
NQ TriviaQA Hellaswag GSM8K MATH MBPP HumanEval
Raw 9.64 59.90 71.56 85.60 23.60 22.40 73.17 49.41
Rand (8K) 12.24 59.22 71.71 83.00 25.2 12.6 73.17 48.16
Inter-Diversity (0-20) 12.94 60.28 71.79 86.20 23.80 11.80 75.61 48.92
Inter-Diversity (40-60) 12.71 59.76 71.71 83.20 28.40 14.40 73.78 49.14
Inter-Diversity (80-100) 11.36 59.52 71.61 83.00 25.40 14.20 68.29 47.63
Intra-Diversity (0-20) 12.33 59.92 71.58 84.00 28.40 19.40 71.34 49.57
Intra-Diversity (40-60) 13.10 59.42 71.64 82.40 21.00 13.20 72.56 47.62
Intra-Diversity (80-100) 12.58 60.06 71.88 84.60 34.90 9.40 73.17 49.51
DaaR Training on Qwen3
Subsequently, we employed DaaR to conduct diversity probe training on Qwen3. Consistent with observations in other models, the two-stage MLP training framework demonstrated robust stability and convergence, shown in Fig. 9.
Figure 9: Training loss and validation process of the two training stages of DaaR on Qwen3-8B, showing the model gradually converging.
Evaluation Results of DaaR with Baselines
Following the identical experimental setup outlined in Appendix B, we conducted experiments on Qwen3, with the results presented in Table 9. As demonstrated, DaaR achieved the top comprehensive score on Qwen3, surpassing all baselines by up to 4.7%. Notably, its performance on the MATH benchmark was particularly outstanding, exceeding the average of baselines by 25.91%. The comparative experiments affirm that DaaR maintains competitive effectiveness even on the SOTA LLM Qwen3-8B.
Table 9: Performance of DaaR with baselines on Qwen3-8B across various benchmarks.
Qwen3-8B Common Sense Reasoning Mathematics Coding Avg
NQ TriviaQA Hellaswag GSM8K MATH MBPP HumanEval
Raw 9.64 59.90 71.56 85.60 23.60 22.40 73.17 49.41
Rand (8K) 12.24 59.22 71.71 83.00 25.20 12.60 73.17 48.16
Instruction Len 11.66 59.56 71.67 83.80 26.30 18.20 73.17 49.19
Alpagasus [10] 12.52 59.34 71.69 83.80 26.60 11.20 69.51 47.81
Instag-C [30] 11.91 59.62 71.55 83.00 26.40 14.60 73.78 48.69
Instag-D [30] 11.99 59.48 71.59 87.20 25.30 13.60 71.34 48.64
SuperFilter [26] 15.73 60.18 71.69 84.80 30.90 16.20 67.07 49.51
Deita-C [29] 12.55 59.62 71.78 83.60 28.60 13.20 77.44 49.54
Deita-Q [29] 12.88 59.72 71.64 85.40 32.90 13.00 73.78 49.90
Deita-D [29] 13.88 59.98 71.50 83.60 26.30 14.00 70.73 48.57
DaaR (Ours) 13.68 60.14 71.56 83.40 34.50 14.60 72.56 50.06
F.3 Customized Data Selection Capability of DaaR
A distinctive characteristic of our method stems from its inherent capability to perform domain-specific data selection under a domain-undetermined scenario, thereby naturally supporting customized data curation. In practical scenarios beyond comprehensive LLM improvement, mission-critical applications (e.g., developing specialized LLM variants for coding or mathematical reasoning) often demand focused enhancement on target capabilities. Conventional baseline selection approaches, which lack domain-aware mechanisms, fundamentally preclude such fine-grained control.
To validate this customized selection capability, we conduct a validated experiment on Qwen2.5-7B. By strategically allocating higher selection ratios (50%) to specific target domains while reducing others to 10%, while maintaining the total training corpus size at 8,000 instances, we investigate whether such curated data distribution induces domain-specific performance gains.
As presented in Table 10, our analysis reveals that compared to the strongest baselines (Random and original DaaR), customized DaaR achieves SOTA enhancement in focused domains, despite observing marginal performance degradation in non-target domains that leads to slightly inferior overall performance compared to the original DaaR. Interestingly, the performance on Reasoning-Major takes first place on Avg score, highlighting that the reasoning capability may be the crucial factor of LLMs. This validation promisingly confirms the operational feasibility of our customization paradigm and demonstrates DaaR’ unique superiority over existing baselines in enabling domain-prioritized LLM specialization.
Table 10: Validation results of customized selected ratio on Qwen2.5-7B across benchmarks.
Qwen2.5-7B Common Sense Reasoning Mathematics Coding Avg
NQ TriviaQA Hellaswag GSM8K MATH MBPP HumanEval
Raw 8.84 58.14 72.75 78.20 9.10 7.40 78.05 44.64
Rand (8K) 11.46 57.85 73.08 78.90 13.15 62.50 71.65 52.66
DaaR (Original) 15.83 58.65 72.48 80.20 16.70 64.20 68.29 53.76
Common-Major 17.56 59.58 72.51 77.80 12.85 62.10 67.53 52.85
Reasoning-Major 16.56 58.22 74.62 77.30 17.20 63.40 66.42 53.39
Math-Major 13.20 57.96 74.16 81.40 16.05 61.80 63.45 52.75
Coding-Major 10.76 57.02 72.78 78.20 12.80 65.20 73.22 52.85
F.4 Rationale for Entropy over Diversity as the Reward Signal
As discussed in Section 4, the explicit diversity metric defined in Eq. (2) & (3) inherently depends on cross-sample relationships that cannot be captured through conventional sample-level training paradigms. To empirically validate this limitation, we conducted an ablation study using 500 representative samples. The inter-diversity scores were calculated using Eq. (2) and partitioned into ascending quartiles, each representing distinct diversity levels. A 5-layer MLP classifier was subsequently trained with cross-entropy loss as a more tractable verification approach compared to regression. Experimental result (Fig. 10) demonstrates persistent convergence challenges, evidenced by consistently high training loss and subpar classification accuracy capped at 34%. These findings substantiate our hypothesis that LLMs fail to learn cross-sample characteristics, thereby necessitating entropy-based proxy metrics for effective reward formulation.
Figure 10: Validation loss and accuracy on directly use diversity instead of entropy as a reward signal.
Appendix G Complete Ablation Studies and Discussion
G.1 Similarity in Generated Domains’ Centroid
G.2 DaaR Layer Selection Protocol
We choose layer-3 as the hidden layer attached to $\psi_{dom}$. To determine an appropriate layer for embedding DaaR into an LLM that balances performance and computational cost, we first evaluate the capacity of different layers for domain awareness. Using Qwen2-7B as an example, we conduct t-SNE visualizations of Layer-5, Layer-10, Layer-15, Layer-20, Layer-25, and the last hidden layer, based on the average token vectors, as shown in Fig. 12.
The observations demonstrate that LLMs maintain cross-dataset classification capabilities across all architectural depths. Considering the diminishing marginal utility of computational resources in deeper layers, we prioritize initial layers to optimize learning efficiency. The training dynamics across the first ten layers are visualized in Fig. 13, demonstrating consistent accuracy within the 91.2–93.6% range, verifying that various layers can effectively support DaaR first-stage training.
Figure 13: Validation loss and accuracy over steps on different layers.
Given that data clustering is based on the embedding layer, a natural approach was to directly connect to the embedding layer for training. However, the training results showed persistently high training loss and a validation accuracy of only about 0.6 (shown in Fig. 14), indicating suboptimal learning performance. Furthermore, related studies indicate that LLMs can learn semantic understanding and perception in the early layers [40], consequently, we select Layer-3 for DaaR.
Figure 14: Training loss and validation process of Embedding Layer on Qwen2.5-7B.
G.3 Ablation of the Number of Seed Samples
To investigate the impact of the number of seed samples $\mathcal{S}_{k}^{(0)}$ on domain centroid construction, we systematically selected 3, 5, and 10 as the number of seed samples while maintaining all other parameters at their default settings on Qwen2.5-7B. For each configuration, pairwise cosine similarity between domains was computed to assess whether significant discrepancies emerged across different seed sizes. Experimental results summarized in Table 11, demonstrate that similarities remain highly consistent across all domain pairs. Except for the common sense domain (with similarity approaching 0.9), varying seed sample counts do not yield substantial differences in the final centroids, with variance ranging from 5.4e-7 to 1.8e-5, demonstrating the stability of this hyperparameter.
Intuitively, we posit that excessively small seed sizes may compromise domain characterization due to generation randomness, while larger sizes risk introducing redundant patterns. We conjecture that the relatively lower absolute similarity in common sense domains stems from inherent sentence brevity, which increases vulnerability to lexical biases. However, this does not impair the anchoring effectiveness, as evidenced by the consistent performance across configurations.
Table 11: Semantic similarity between the number of seed samples among 3, 5 and 10 across domains on Qwen2.5-7B.
Domains Similarity of 3 & 5 Similarity of 3 & 10 Similarity of 5 & 10
Common Sense 0.8984 0.8913 0.9074
Reasoning 0.9713 0.9752 0.9815
Mathematics 0.9776 0.9761 0.9722
Coding 0.9775 0.9759 0.9760
G.4 Ablation of the Number of Diversity Augmentation Data
To assess the impact of the amount of generated data on the accuracy of domain semantic centroids, we select quantities of 10 and 30 samples for generation. Subsequently, we calculate the semantic cosine similarity between these two sets, with results presented in Table 12 (using Qwen2.5-7B and Llama3.1-8B as examples). The observed differences between varying data quantities were not significant, indicating that the generalization of the generated data is sufficient, and further increasing the data quantity does not significantly enhance the accuracy of the centroids.
Table 12: Semantic similarity between generated samples of size 10 and 30 across different domains.
Similarity of 10 & 30 Common Sense Reasoning Mathematics Coding
Qwen2.5-7B 0.9686 0.9866 0.9919 0.9881
Llama3.1-8B 0.9851 0.9795 0.9685 0.9895
G.5 Ablation of the Size of Sliding Window
To investigate the impact of the sliding window size in diversity augmentation, we kept all other hyper-parameters constant, setting the sliding window size to 1, 3, and 5 examples, respectively, and compared the similarity between them. The experimental results on Qwen2.5-7B, presented in Table 13, indicate that different window sizes maintain a high degree of similarity, suggesting that DaaR does not heavily depend on this hyperparameter choice. Considering the prompt length and stability, we opted for a window size of 3 in the implementation.
Table 13: Semantic similarity between the size of sliding window among 1, 3 and 5 across domains on Qwen2.5-7B.
Domains Similarity of 1 & 3 Similarity of 1 & 5 Similarity of 3 & 5
Common Sense 0.9205 0.9094 0.9022
Reasoning 0.9746 0.9760 0.9772
Mathematics 0.9736 0.9618 0.9847
Coding 0.9849 0.9647 0.9716
G.6 Ablation of the Diversity Threshold
The threshold $\tau$ design needs to satisfy two objectives: (1) mitigate redundancy in LLM-generated outputs by preventing exact duplicates in synthetic data, and (2) ensure sufficient similarity among instances within the same domain embedding space. As illustrated in Table 14, we conduct an experiment using different similarity thresholds across various domains on Qwen2.5-7B, observing the number of attempts required for successful generation. The results indicate that distinct domains exhibit varying sensitivity to similarity metrics, with the number of attempts increasing sharply as the threshold decreases. Based on these observations, we empirically determined a threshold value of 0.9 for Mathematics and 0.85 for others as a balanced compromise.
Table 14: Generation attempt frequencies for 10 seed samples across different domains and similarity levels on Qwen2.5-7B.
Domains 0.8 0.85 0.9 0.95
Common Sense 153 31 21 12
Reasoning 135 61 44 15
Mathematics $>1000$ 804 236 25
Coding 74 21 16 13
We further conducted ablation experiments by comparing the similarity of embedding centroids generated under different thresholds to systematically evaluate the impact of threshold variations on downstream performance. The results presented in Table 15 demonstrate, with the exception of common sense data (similar to the results in Table 11), that varying thresholds do not significantly affect the role of these centroids as initial points for clustering. We conjecture that lower thresholds ensure greater diversity within domain, whereas higher thresholds tend to result in repetitive data.
Table 15: Semantic similarity between the similarity threshold among 0.85, 0.9, and 0.95 across domains on Qwen2.5-7B.
Domains Similarity of 0.85 & 0.9 Similarity of 0.85 & 0.95 Similarity of 0.85 & 0.95
Common Sense 0.8869 0.8963 0.9502
Reasoning 0.9791 0.9785 0.9795
Mathematics 0.9730 0.9754 0.9703
Coding 0.9646 0.9740 0.9691
G.7 Cost-Efficiency and Flexibility
Compared to baseline methods requiring GPT-based evaluators (Alpagasus, Instag) or full-LLM inference (Deita), our approach achieves superior efficiency through data-model co-optimizations. Our method demonstrates computational efficiency through two key aspects enabled by the LLM’s self-rewarding capability during dedicated data synthesis:
  • Our framework operates on frozen embeddings extracted from layer 3 while truncating subsequent layers, which constitutes a significantly shallower architecture compared to the conventional 32-layer structure in 7B-scale LLMs. Taking Qwen2 as an example, full model inference requires loading complete parameters alongside cache management components, consuming 18–24 GB GPU memory. In contrast, our method accomplishes data filtering with merely 5–6 GB GPU memory footprint.
  • The lightweight 5-layer MLP probe module introduces only 76 million additional parameters. As a regression model producing single scalar outputs per instance, it achieves substantial speed improvements over LLM-inference-based approaches. Experimental results on 5,000 samples demonstrate this efficiency advantage: our method requires approximately 40 minutes for sample-level inference, whereas conventional LLM inference approaches demand nearly two hours.
G.8 Stability of DaaR
Our dual-MLP design of DaaR strategically separates entropy regression (second MLP) from neural softmax-based measurement (first MLP), both showing stable convergence in Fig 2. While precise analysis of compounding errors remains a theoretical concern, our framework inherently mitigates these errors through these aspects: (1) Seed Generation: LLM-generated descriptions exhibit minimal variation across iterations, and repeated seeding mitigates uncertainty. (2) Layer Selection: Semantic feature extraction across layers is demonstrated to effectively capture domain-specific patterns, thereby establishing methodological robustness, detailed in Appendix G.2. (3) Clustering: Data seeding as initialization for K-means clustering, where minimal perturbation of initial points ensures stable training of the first MLP. (4) Selection Strategy: Percentage-based data selection prioritizing sample distribution robustness over fine-grained entropy value utilization.
Appendix H Comprehensive Visualization and Experimental Results
H.1 Visualization of Embeddings on Llama3.1-8B , Qwen2-7B & Qwen2.5-7B
We process concatenated data samples with “instruction”+“input”+“output” pairs through each LLM’s embedding layer, computing mean token embeddings for visualization. We present the t-SNE visualization of data samples from Llama3.1-8, Qwen2-7B and Qwen2.5-7B with different distributions in Fig. 15, Fig. 16 and Fig. 17. It is evident that, despite differences in architecture or model, the method of filtering data through Inter-Diversity and Intra-Diversity is effective. Notably, although the embedding distributions of different models are not identical, they exhibit similar behavior, indicating that the representations learned by the embedding layers are comparable.
Figure 15: The t-SNE visualization of embeddings for data samples with different distributions on Llama3.1-8B. (a) The data pool of all 40K samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids on Inter-Diversity, (d) Distribution of data closest to other domain centroids on Inter-Diversity, (e) Distribution of data closest to its own domain centroid on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid on Inter-Diversity.
Figure 16: The t-SNE visualization of embeddings for data samples with different distributions on Qwen2-7B. (a) The data pool of all 40K samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids on Inter-Diversity, (d) Distribution of data closest to other domain centroids on Inter-Diversity, (e) Distribution of data closest to its own domain centroid on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid on Inter-Diversity.
Figure 17: The t-SNE visualization of embeddings for data samples with different distributions on Qwen2.5-7B. (a) The data pool of all 40K samples, (b) Randomly selected subset, (c) Distribution of data farthest from other domain centroids on Inter-Diversity, (d) Distribution of data closest to other domain centroids on Inter-Diversity, (e) Distribution of data closest to its own domain centroid on Inter-Diversity, (f) Distribution of data farthest from its own domain centroid on Inter-Diversity.
H.2 Visualization of Inter- & Intra-Diversity Distribution Data
@@
H.3 Comprehensive Training Dynamics
@@
H.4 Complete Validation Results of Inter-Diversity and Intra-Diversity
@@
Table 16: Validation results of Inter-Diversity and Intra-Diversity on Llama3.1-8B across benchmarks.
@@
Table 17: Validation results of Inter-Diversity and Intra-Diversity on Qwen2-7B across benchmarks.
@@
Table 18: Validation results of Inter-Diversity and Intra-Diversity on Qwen2.5-7B across benchmarks.
@@
H.5 Complete Results of DaaR with Baselines
@@
Alpagasus [10] 21.57 64.37 74.87 55.20 17.65 4.60 16.16 36.34
Instag-C [30] 18.12 64.96 74.01 55.70 15.50 4.80 37.81 38.70
Instag-D [30] 21.94 64.69 74.87 54.80 12.80 4.10 9.76 34.71
SuperFilter [26] 22.95 64.99 76.39 57.60 6.05 2.60 40.55 38.73
Deita-C [29] 15.58 64.97 74.21 55.00 13.05 4.60 34.46 37.41
Deita-Q [29] 19.57 64.22 75.15 54.00 7.20 4.20 28.35 36.10
Deita-D [29] 20.97 63.32 75.10 54.90 7.00 4.00 31.71 36.71
@@
Alpagasus [10] 13.67 57.94 73.04 73.90 32.30 51.40 63.41 52.24
Instag-C [30] 9.51 58.50 73.06 74.70 35.35 51.90 64.70 52.53
Instag-D [30] 12.87 57.48 72.80 74.40 33.75 51.80 64.02 52.45
SuperFilter [26] 19.16 58.98 72.99 73.70 30.10 52.40 58.85 52.31
Deita-C [29] 8.94 58.07 73.06 73.90 35.55 52.90 62.20 52.09
Deita-Q [29] 14.06 59.07 73.16 75.80 35.50 23.00 58.24 48.40
Deita-D [29] 16.41 57.80 72.70 76.10 29.05 52.40 64.63 52.73
@@
Alpagasus [10] 10.40 57.87 72.92 77.20 18.75 61.80 65.55 52.07
Instag-C [30] 10.81 58.45 73.27 76.00 13.30 61.80 68.29 51.70
Instag-D [30] 11.08 58.40 72.79 76.40 16.40 62.90 70.43 52.63
SuperFilter [26] 13.54 58.51 72.89 79.30 11.35 39.50 65.25 48.62
Deita-C [29] 10.50 58.17 73.14 74.60 16.60 62.00 72.26 52.47
Deita-Q [29] 11.24 57.83 72.97 78.50 12.95 38.10 67.68 48.47
Deita-D [29] 10.48 57.81 73.05 77.20 15.25 52.90 69.21 50.84